# San Jose/Seattle, US https://bytedance.github.io/vidi-website/

### v2

@@ 1 Introduction
Video has become one of the most dominant mediums for sharing information on the Internet. However, for most users, video creation remains a complex and time-consuming process, especially on mobile devices where precise editing is challenging. Among all editing stages, the most labor-intensive step is often identifying the desired segments with long, unedited footage. Beyond trimming, users frequently struggle with video composition tasks such as selecting appropriate music, transitions, effects, animations, filters, stickers, and fonts, which require both technical skill and artistic judgment. Moreover, creative editing actions beyond composition are increasingly important in modern video production but are even harder to achieve without expert tools or AI assistance, e.g., generating thumbnails, cover art or stylized scenes. We aim to build the next generation of video creation systems to be powered by advanced video understanding and editing capabilities, enabling users to complete complex editing workflows automatically and effortlessly.
Figure 2: Examples of temporal retrieval queries and their corresponding time ranges from the proposed VUE-TR benchmark. The duration of the example video is $3,871$ seconds (i.e., 01:04:31). Each query is presented in one of three formats (i.e., keyword, phrase, sentence) with visual or audio information. Facial regions in the video frames are blurred to preserve privacy.
@@
To support comprehensive evaluation in realistic settings, we introduce a manually annotated benchmark (abbreviated as VUE-TR) for temporal retrieval. VUE-TR consists of videos ranging from approximately $20$ seconds to over $1$ hour, categorized into five groups: ultra-short ($<1$ min), short ($1-10$ mins), medium ($10-30$ mins), long ($30-60$ mins) and ultra-long ($>1$ hour). Each video is paired with queries of varying formats and lengths (i.e., keyword, phrase, and sentence) to reflect the diversity of user search intent. Critically, the queries may require visual, audio, or both modalities for accurate localization. This aligns closely with real-world scenarios, where audio plays an essential role in video comprehension, particularly in domains such as TV shows, broadcast, and musical performances. As illustrated in Figure 2, we showcase an example from an hour-long video with three representative queries, spanning different formats and modalities. Remarkably, Vidi-7B significantly outperforms leading LMMs (Large Multimodal Models), e.g., Gemini and GPT-4o, highlighting its effectiveness in complex, multimodal temporal retrieval tasks.
@@ 2 General Overview
Figure 3 illustrates the model architecture of Vidi. For LLM training, we adopt the decomposed attention mechanism proposed by Kuo et al. [13], which reduces the computation complexity over $N$ multimodal tokens from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ without sacrificing performance on downstream multimodal tasks. This efficient attention design enables both training and inference on extremely long videos, which are otherwise infeasible for standard transformer-based models [27]. All videos are uniformly sampled at $1$ frames per second (fps) for visual input and $16,000$ Hz for audio, ensuring that the model can localize and understand content at second-level precision.
@@
The training of Vidi is conducted in two main stages to ensure strong multimodal temporal retrieval:
@@
  • Application Post-training: To support diverse application scenarios, we further fine-tune the model on task-specific data, including temporal retrieval and video/image VQA. The temporal retrieval training data mirrors the structure of the VUE-TR benchmark, with queries in three formats (i.e., keyword, phrase, sentence), requiring information from vision, audio, or both. This stage enhances the model’s ability to handle real-word usage patterns where query types vary significantly.
Depending on the intended use case, we train different versions of Vidi, with the tailored training recipes, e.g., a temporal retrieval expert model, and a more generic VQA model (to be included in future releases). During inference, Vidi is capable of running on a single 80G GPU without quantization, and efficiently process videos exceeding $2$ hours in length, making it practical for real-world deployment in long-form video editing and retrieval tasks.
@@ 3 Model Architecture
Unlike generic video understanding tasks that rely on sparse or uniform sampling of a fixed number of frames, temporal retrieval requires models to localize relevant segments with second-level precision, necessitating dense sampling, typically at $1$ fps. Under this fixed sampling rate, the multimodal inputs (e.g., vision, audio, and text) can vary significantly in length, ranging from ultra-short of just a few seconds to ultra-long exceeding an hour. Furthermore, our proposed VUE-TR benchmark emphasizes realistic retrieval conditions by requiring models to process both visual and audio modalities when localizing queries. To address challenges in handling videos with varying length, multimodal inputs at scale, we adopt the Decomposed Attention (D-Attn[13] architecture in Vidi. D-Attn offers exceptional computational efficiency while maintaining strong multimodal capabilities. It also facilitates seamless integration of visual, audio, and textual streams, making it well-suited for temporal retrieval in long-form, real-world videos.
As shown in Figure 3, the visual and audio inputs are first encoded into token sequences using the pretrained visual and audio encoders, respectively. The resulting multimodal embeddings are projected with corresponding adapter layers. Together with the text query embeddings, they are fed into a D-Attn LLM, which localizes the text query within the input video.
@@
From a computational standpoint, a diagonalized variant of V2V Self-Attn is proposed to retain performance while reducing complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ for $N$ visual tokens, shown in Figure 4. This lightweight design is particularly well-suited for densely sampled, ultra-long videos, where $N$ might be extremely large. For example, if each frame consists of 400 tokens, a one-hour video would have 1.44M tokens. Such input length is challenging for conventional fully self-attention mechanisms.
@@
To effectively address the challenge of varying-length video input, we modify the $\alpha$-weighting strategy used in the D-Attn framework. This adjustment is designed to balance the contributions of textual and multimodal information. In particular, video lengths (and thus token counts) can vary significantly in the context of temporal retrieval. In D-Attn, Kuo et al. [13] analytically derive how self-attention can be equivalently decomposed into a weighted sum of T2T Self-Attn and T2V Cross-Attn. Given a text token $t$ and a sequence of concatenated visual and textual tokens $[V,T]$, where $V=[v_{1},v_{2},\cdots,v_{N}]$ represents a sequence of visual tokens and $T=[t_{1},t_{2},\cdots,t_{M}]$ represents a sequence of text tokens, the attention from $t$ to $[V,T]$ can be formulated as
@@
where $\alpha_{V}=\texttt{Sigmoid}(S_{V}-S_{T})$, $\alpha_{T}=\texttt{Sigmoid}(S_{T}-S_{V})=1-\alpha_{V}$. $S_{V}$ and $S_{T}$ are defined as the log sum of exponential of the dot product between $t$ and $V$, and between $t$ and $T$, respectively.
@@
where $N$ is the number of visual tokens and $M$ is the number of text tokens. In the temporal retrieval task, this formulation leads to an imbalance: while the text query length $M$ typically stays within a small range, the visual token count $N$ can vary dramatically due to differences in video duration. As a result, $S_{V}$ tends to be significantly larger for longer videos, which in turn biases $\alpha_{V}$ towards $1$ and $\alpha_{T}$ towards 00. This causes the model to overemphasize visual inputs while neglecting the textual query, potentially degrading performance. To mitigate this issue, we simplify the formulation in Vidi by fixing the weighting coefficients, i.e., $\alpha_{V}=\alpha_{T}=1$. This ensures a balanced contribution from both modalities, regardless of the input video length.
The D-Attn framework can be trivially generalized to accommodate audio inputs. Given audio tokens $A=[a_{1},a_{2},\cdots,a_{P}]$, the attention from a text token $t$ to the combined multimodal sequence $[V,A,T]$ can be similarly derived as
@@ 4.1 Adapter Training
In this stage, we train the adapters from scratch while keeping the weight of the vision/audio encoders and the LLM fixed. Each adapter contains a convolution layer for compressing the raw visual or audio tokens, followed by an MLP that maps the compressed representations into the LLM-compatible input space. We leverage the strong publicly available pretrained models for all other components: SigLIP [34] for vision, Whisper [22] for audio, and Mistral [11] for the language model backbone. The adapters are trained on approximately $1$ million image and audio caption data, allowing them to learn semantic grounding across modalities while maintaining alignment with the LLM.
@@ 4.2 Synthetic Data Training
To support large-scale training for temporal alignment, we propose a synthetic video generation pipeline based on captioned video and audio datasets with ground-truth timestamps. As shown in Figure 5, we begin by randomly sampling a group of candidate samples from large-scale caption corpora, i.e., approximately $25$ million images and $400$ thousand audio clips. Each image is expanded into a synthetic video segment using a sliding window approach, with randomized parameters such as window size, starting corner, sliding direction, and sliding speed. This mimics camera movement or visual variation over time. Similarly, audio clips are segmented and spliced in random order to create synthetic audio tracks. The resulting synthetic inputs are arbitrarily long and composed of diverse multimodal segments. Since each segment originates from a known captioned sample, we can automatically generate large-scale (timestamp, caption) pairs as supervision. Since images do not match audio, the training input contains only one active modality (vision or audio), while the other is padded with empty tokens.
We train the model on two complementary objectives: 1) caption prediction: given a timestange range, predict the corresponding caption; 2) timestamp localization: given a caption, predict the matching timestamp range. These dual tasks help the model develop a deep understanding of both visual/audio content and the temporal axis. After training on synthetic data, the model achieves over $80\%$ accuracy on a synthetic evaluation set with multi-minute sequences, indicating strong alignment capability and robust temporal grounding.
@@ 4.3 Real Video Training
To bridge the gap between synthetic videos and real-world videos, we further train the model on a large corpus of $1$ million long videos, each annotated with dense supervision in the form of (timestamp, caption) and (timestamp, subtitle) pairs. As shown in Figure 6, each long video is first segmented into short clips (typically $5-30s$) using a combination of scene boundary detection [16] and subtitle punctuation cues [8]. This process yields approximately $3-500$ segments per video, depending on content density and duration. After that, dense captions are generated with state-of-the-art LMMs [3, 36], resulting in fine-grained semantic coverage. Original subtitles are reprocessed into sentence-level units using punctuation-based splitting, enhancing their suitability for timestamped supervision. This pipeline produces over $30$ million paired (timestamp, caption) and (timestamp, subtitle) training samples. The training input consists of raw video files containing both visual and audio tracks.
@@
We design four training objectives to fully leverage the aligned supervision:
@@
These tasks not only reinforce the model’s ability to align textual descriptions with multimodal signals but also help it adapt to complex video/audio in real-world distributions.
@@ 5 Application Post-training
  • sentence: complete sentence(s) describing detailed events or scenes, e.g., “The majestic presence of a volcano surrounded by lush vegetation and shrouded by clouds”.
@@
We design a rule-based post-processing step to enhance the quality of generated queries and the corresponding time ranges, formed by the following stages. 1. Merging adjacent time ranges: for each query, we combine consecutive time segments or those separated by small gaps (e.g., $0.5$ seconds), provided the query appearing in captions or subtitles. 2. Filtering out low-confidence queries: queries with confidence scores below $0.9$ are discarded to ensure reliability. 3. Excluding overly general queries: queries associated with more than $10$ timestamps are considered too general. We remove such cases that frequently appear throughout the video. 4. Eliminating machine-style queries: queries exhibiting templated or unnatural phrasing, such as “The video concludes…” or “In the closing moments,” will be filtered out to ensure quality. After that, the generated queries and the corresponding time ranges are sent to human annotators to conduct the final stage verification and modification.
@@
Although involving a filtering process, generated queries and the corresponding time ranges still contain ambiguity of phrasing and misaligned timestamps. To obtain high-quality annotations, we further design a comprehensive human-in-the-loop refinement process. Annotators begin by examining each query for clarity and relevance. If a query is ambiguous or unrealistic, it is required to be rewritten to reflect the video content accurately. After that, annotators review the entire video to identify precise time segments corresponding to the refined queries. Besides, another annotator is required to review annotations independently. Discrepancies identified during this phase will be resolved through discussion or by consulting a senior annotator.
@@
  • Vision. The query can be accurately identified purely based on the visual content. Audio information is not necessary for retrieving the segments.
  • Audio. The query depends solely on auditory information. Visual cues are not necessary to identify the video segment.
  • Vision+Audio. The query relies on both visual and auditory information for accurate interpretation and localization within the video.
This categorization strategy helps researchers to evaluate the capabilities of the models with a more detailed analysis.
@@ 6.1 Data Statistics
  • Video Duration. Unlike prior datasets limited to short clips, the video length varies from $20$ seconds to over $1$ hour, covering the full spectrum of durations encountered in real-world scenarios. To facilitate duration-wise evaluation, we categorize videos into five balanced buckets: ultra-short ($<1$ min), short ($1-10$ mins), medium ($10-30$ mins), long ($30-60$ mins) and ultra-long ($>60$ mins). It enables systematic analysis of model performance as video length increases, which is a key challenge for temporal retrieval that prior works fail to capture.
  • Query Modality. While most large vision-language models (LVLMs) operate on vision and text alone, VUE-TR explicitly integrates audio as a core input modality. As shown in Figure 8 (left), $47\%$ of the queries involve both visual and audio signals, $35\%$ are vision-only, and $18\%$ are audio-only. This breakdown enables comprehensive evaluation of multimodal capabilities.
  • Query Format. VUE-TR is the first temporal retrieval benchmark to incorporate multiple query formats, reflecting the diversity of user intent in real-word search scenarios. The three formats, ranging from keywords to free-form sentences, are carefully balanced, as shown in Figure 8 (right). Then the model can be evaluated to handle queries of varying linguistic complexity.

### v3

@@ 1 Introduction
Video has become one of the most dominant mediums for sharing information on the Internet. However, for most users, video creation remains a complex and time-consuming process, especially on mobile devices where precise editing is challenging. Among all editing stages, the most laborious step is often identifying the desired segments with long, unedited footage. Beyond trimming, users frequently struggle with video composition tasks, such as selecting appropriate music, transitions, effects, animations, filters, stickers, and fonts, which require both technical skill and artistic judgment. Moreover, creative editing actions beyond composition are increasingly important in modern video production but are even harder to achieve without expert tools or AI assistance, e.g., generating thumbnails, cover art or stylized scenes. We aim to build the next generation of video creation systems powered by advanced video understanding and editing capabilities, enabling users to complete complex editing workflows automatically and effortlessly.
Figure 2: Examples of temporal retrieval queries and their corresponding time ranges from the proposed VUE-TR benchmark. The duration of the example video is $3,871$ seconds (i.e., 01:04:31). Each query is presented in one of three formats (i.e., keyword, phrase, sentence) with visual or audio information. Facial regions in the video frames are blurred to protect privacy.
@@
To support comprehensive evaluation in realistic settings, we introduce a manually annotated benchmark (abbreviated VUE-TR) for temporal retrieval. VUE-TR consists of videos ranging from approximately $20$ seconds to over $1$ hour, categorized into five groups: ultra-short ($<1$ min), short ($1-10$ mins), medium ($10-30$ mins), long ($30-60$ mins) and ultra-long ($>1$ hour). Each video is paired with queries of varying formats and lengths (i.e., keyword, phrase, and sentence) to reflect the diversity of the user search intent. Critically, queries may require visual, audio, or both modalities for accurate localization. This aligns closely with real-world scenarios, where audio plays an essential role in video comprehension, particularly in domains such as TV shows, broadcasts, and musical performances. As illustrated in Figure 2, we present an example of an hour-long video with three queries, covering different formats and modalities. Remarkably, Vidi-7B significantly surpasses leading Large Multimodal Models (LMMs), such as Gemini and GPT-4o, highlighting its effectiveness in handling complex multimodal temporal retrieval tasks. Building upon this success, Vidi-1.5-9B achieves further performance gains through an enhanced LLM backbone, an upgraded visual encoder, an adaptive token compression strategy, and the incorporation of additional high-quality training data.
@@ 2 General Overview
Figure 3 illustrates the model architecture of Vidi. For LLM training, we adopt the decomposed attention mechanism proposed by Kuo et al. [13], which reduces the computational complexity over multimodal tokens $N$ from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ without sacrificing performance on downstream multimodal tasks. This efficient attention design enables training and inference in extremely long videos, which are otherwise infeasible for standard transformer-based models [28]. All videos are uniformly sampled at $1$ frames per second (fps) for visual input and $16,000$ Hz for audio, ensuring that the model can localize and understand content with second-level precision.
@@
Training of Vidi is carried out in two main stages to ensure strong multimodal temporal retrieval:
@@
  • Application Post-training: To support diverse application scenarios, we further fine-tune the model on task-specific data, including temporal retrieval and video/image VQA. The temporal retrieval training data mirror the structure of the VUE-TR benchmark, with queries in three formats (i.e., keyword, phrase, sentence), requiring information from vision, audio, or both. This stage enhances the model’s ability to handle real-word usage patterns where query types vary significantly.
Depending on the intended use case, we train different versions of Vidi, with custom training recipes, e.g., a temporal retrieval expert model, and a more generic VQA model (to be included in future releases). During inference, Vidi is capable of running on a single 80G GPU without quantization, and efficiently process videos exceeding $2$ hours in length, making it practical for real-world deployment in long-form video editing and retrieval tasks.
@@ 3 Model Architecture
Unlike generic video understanding tasks that rely on sparse or uniform sampling of a fixed number of frames, temporal retrieval requires models to localize relevant segments with second-level precision, necessitating dense sampling, typically at $1$ fps. Under this fixed sampling rate, the multimodal inputs (e.g., vision, audio, and text) can vary significantly in length, ranging from ultra-short of just a few seconds to ultra-long exceeding an hour. Furthermore, our proposed VUE-TR benchmark emphasizes realistic retrieval conditions by requiring models to process both visual and audio modalities when localizing queries. To address challenges in handling videos with varying length and multimodal inputs at scale, we adopt the Decomposed Attention (D-Attn[13] architecture in Vidi. D-Attn offers exceptional computational efficiency while maintaining strong multimodal capabilities. It also facilitates seamless integration of visual, audio, and textual streams, making it well suited for temporal retrieval in long-form, real-world videos.
As shown in Figure 3, the visual and audio inputs are first encoded into token sequences using pretrained visual and audio encoders, respectively. The resulting multimodal embeddings are projected with corresponding adapter layers. Together with the text query embeddings, they are fed into a D-Attn LLM, which localizes the text query within the input video.
@@
From a computational point of view, a diagonal variant of V2V Self-Attn is proposed to retain performance while reducing the complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$ for the visual tokens $N$, shown in Figure 4. This lightweight design is particularly well suited for densely sampled, ultra-long videos, where $N$ might be extremely large. For example, if each frame consists of 400 tokens, a one-hour video would have 1.44M tokens. Such input length is challenging for conventional fully self-attention mechanisms.
@@
To effectively address the challenge of varying-length video input, we modify the $\alpha$-weighting strategy used in the D-Attn framework. This adjustment is designed to balance the contributions of textual and multimodal information. In particular, video lengths (and thus token counts) can vary significantly in the context of temporal retrieval. In D-Attn, Kuo et al. [13] analytically derive how self-attention can be equivalently decomposed into a weighted sum of T2T Self-Attn and T2V Cross-Attn. Given a text token $t$ and a sequence of concatenated visual and textual tokens $[V,T]$, the attention from $t$ to $[V,T]$ can be formulated as
@@
where $\alpha_{V}=\texttt{Sigmoid}(S_{V}-S_{T})$, $\alpha_{T}=\texttt{Sigmoid}(S_{T}-S_{V})=1-\alpha_{V}$. $V=[v_{1},v_{2},\cdots,v_{N}]$ and $T=[t_{1},t_{2},\cdots,t_{M}]$ represent a sequence of visual and text tokens, respectively. $S_{V}$ and $S_{T}$ are defined as the logarithmic sum of the exponential of the dot product between $t$ and $V$, and between $t$ and $T$, respectively.
@@
where $N$ is the number of visual tokens and $M$ is the number of text tokens. In the temporal retrieval task, this formulation leads to an imbalance: while the length of the text query $M$ typically stays within a small range, the count of visual tokens $N$ can vary dramatically due to differences in the duration of the video. As a result, $S_{V}$ tends to be significantly larger for longer videos, which in turn biases $\alpha_{V}$ towards $1$ and $\alpha_{T}$ towards 00. This causes the model to overemphasize visual input while neglecting the textual query, potentially degrading performance. To mitigate this issue, we simplify the formulation in Vidi by fixing the weighting coefficients, i.e., $\alpha_{V}=\alpha_{T}=1$. This ensures a balanced contribution from both modalities, regardless of the input video length.
The D-Attn framework can be trivially generalized to accommodate audio inputs. Given audio tokens $A=[a_{1},a_{2},\cdots,a_{P}]$, the attention of a text token $t$ to the combined multimodal sequence $[V,A,T]$ can be similarly derived as
@@ 4.1 Adapter Training
In this stage, we train the adapters from scratch while keeping the weight of the vision/audio encoders and the LLM fixed. Each adapter contains a convolution layer for compressing the raw visual or audio tokens, followed by an MLP that maps the compressed representations into the LLM-compatible input space. We leverage the strong publicly available pretrained models for all other components: SigLIP series [35, 27] for vision, Whisper [22] for audio, and Mistral [11]/Gemma [25] for the language model backbone. The adapters are trained on approximately $1$ million image and audio caption data, allowing them to learn semantic grounding across modalities while maintaining alignment with the LLM.
@@ 4.2 Synthetic Data Training
To support large-scale training for temporal alignment, we propose a synthetic video generation pipeline based on captioned video and audio datasets with ground-truth timestamps. As shown in Figure 5, we begin by randomly sampling a group of candidate samples from large-scale caption corpora, i.e., approximately $25$ million images and $400$ thousand audio clips. Each image is expanded into a synthetic video segment using a sliding window approach with randomized parameters such as window size, starting corner, sliding direction, and sliding speed. This mimics camera movement or visual variation over time. Similarly, audio clips are segmented and spliced in random order to create synthetic audio tracks. The resulting synthetic inputs are arbitrarily long and composed of diverse multimodal segments. Since each segment originates from a known captioned sample, we can automatically generate large-scale (timestamp, caption) pairs as supervision. Since images do not match audio, the training input contains only one active modality (vision or audio), while the other is padded with empty tokens.
We train the model on two complementary objectives: 1) caption prediction: given a timestamp range, predict the corresponding caption; 2) timestamp localization: given a caption, predict the matching timestamp range. These dual tasks help the model develop a deep understanding of both visual/audio content and the temporal axis. After training on synthetic data, the model achieves over $80\%$ accuracy on a synthetic evaluation set with multi-minute sequences, indicating strong alignment capability and robust temporal grounding.
@@ 4.3 Real Video Training
To bridge the gap between synthetic videos and real-world videos, we further train the model on a large corpus of $1$ million long videos, each annotated with dense supervision in the form of pairs (timestamp, caption) and (timestamp, subtitle). As shown in Figure 6, each long video is first segmented into short clips (typically $5-30s$) using a combination of scene boundary detection [16] and subtitle punctuation cues [8]. This process yields approximately $3-500$ segments per video, depending on the density and duration of the content. After that, dense captions are generated with state-of-the-art LMMs [3, 37], resulting in fine-grained semantic coverage. Original subtitles are reprocessed into sentence-level units using punctuation-based splitting, enhancing their suitability for timestamped supervision. This pipeline produces more than $30$ million paired training samples (timestamp, caption) and (timestamp, subtitle). The training input consists of raw video files containing both visual and audio tracks.
@@
We design four training objectives to fully leverage aligned supervision.
@@
These tasks not only reinforce the model’s ability to align textual descriptions with multimodal signals, but also help it adapt to complex video/audio in real-world distributions.
@@ 5 Application Post-training
  • sentence: complete sentence(s) describing detailed events or scenes, e.g., “The majestic presence of a volcano surrounded by lush vegetation and shrouded in clouds”.
@@
We design a rule-based post-processing step to enhance the quality of generated queries and the corresponding time ranges, formed by the following stages. 1. Merging adjacent time ranges: for each query, we combine consecutive time segments or those separated by small gaps (e.g., $0.5$ seconds), provided the query appearing in captions or subtitles. 2. Filtering out low-confidence queries: queries with confidence scores below $0.9$ are discarded to ensure reliability. 3. Excluding overly general queries: queries associated with more than $10$ timestamps are considered too general. We remove such cases that frequently appear throughout the video. 4. Eliminating machine-style queries: queries exhibiting templated or unnatural phrasing, such as “The video concludes…” or “In the closing moments,” will be filtered out to ensure quality. After that, the generated queries and the corresponding time ranges are sent to human annotators to perform the final stage verification and modification.
@@
Although involving a filtering process, generated queries and the corresponding time ranges still contain ambiguity of phrasing and misaligned timestamps. To obtain high-quality annotations, we further design a comprehensive human-in-the-loop refinement process. Annotators begin by examining each query for clarity and relevance. If a query is ambiguous or unrealistic, it is required to be rewritten to reflect the video content accurately. After that, the annotators review the entire video to identify precise time segments corresponding to the refined queries. In addition, another annotator is required to review the annotations independently. Discrepancies identified during this phase will be resolved through discussion or by consulting a senior annotator.
@@
  • Vision. The query can be accurately identified purely on the basis of visual content. Audio information is not necessary to retrieve the segments.
  • Audio. The query depends solely on the auditory information. Visual cues are not necessary to identify the video segment.
  • Vision+Audio. The query relies on visual and auditory information for accurate interpretation and localization within the video.
This categorization strategy helps researchers evaluate the capabilities of the models with a more detailed analysis.
@@ 6.1 Data Statistics
  • Video Duration. Unlike prior datasets limited to short clips, the video length varies from $20$ seconds to more than $1$ hour, covering the full spectrum of durations encountered in real-world scenarios. To facilitate duration-wise evaluation, we categorize videos into five balanced buckets: ultra-short ($<1$ min), short ($1-10$ mins), medium ($10-30$ mins), long ($30-60$ mins) and ultra-long ($>60$ mins). It enables systematic analysis of model performance as the video length increases, which is a key challenge for temporal retrieval that prior work failed to capture.
  • Query Modality. While most large vision-language models (LVLMs) operate on vision and text alone, VUE-TR explicitly integrates audio as a core input modality. As shown in Figure 8 (left), $47\%$ of the queries involve both visual and audio signals, $35\%$ are vision-only, and $18\%$ are audio-only. This breakdown enables a comprehensive evaluation of multimodal capabilities.
  • Query Format. VUE-TR is the first temporal retrieval benchmark to incorporate multiple query formats, reflecting the diversity of user intent in real-word search scenarios. The three formats, ranging from keywords to free form sentences, are carefully balanced, as shown in Figure 8 (right). Then the model can be evaluated to handle queries of varying linguistic complexity.